OpenStudy (anonymous):

If the area of a triangle DEF above is equal to the area of square M (not shown), what is the length of a side of the square?

OpenStudy (welshfella):

where is the triangle?

OpenStudy (anonymous):



OpenStudy (welshfella):

its not possible to evaluate the area of that triangle as we need its height as well as its base

OpenStudy (anonymous):

the height is 10 sorry forgot to put it

OpenStudy (welshfella):

Ok area of triangle = (1/2) * base 8 height = 1/2 * 5 * 10 can you work that out?

OpenStudy (anonymous):

yes 25

OpenStudy (welshfella):

* 1/2 * base * height

OpenStudy (welshfella):

ok so if the length of the side of the square is x then its area is x^2 so x^2 = 25 what is x?

OpenStudy (anonymous):

5

OpenStudy (welshfella):

yep

OpenStudy (anonymous):

thank you for the help and explaining it to me

OpenStudy (welshfella):

yw